# Elastic Full-Waveform Inversion : How the physics of problem improves data-driven techniques?

### v1

@@ 5 Third method (New Variational Autoencoder)
We removed constants that are not relevant. Furthermore, |.||.|| . | denotes the matrix’s determinant. Given these assumptions, the generic VAE objective from 17 can be refined to:
@@ 6.1 Generating random field
To produce random fields for the dataset, we set domain $\Omega=[-1,1]\times[-1,1]$ and divide $\Omega$ to $N=n\times n$ uniform rectangular elements. Using observation and experimentation, we create 28 pairs of $(a_{1},a_{2})$ values in such a way that they cover nearly all situations of earth density. For each pair of $(a_{1},a_{2})$ we have a covariance matrix as 22, and construct 1000 realizations from it. In Figure 8, realizations obtained from specific pairs of $(a_{1},a_{2})$ are observable.
@@ 6.2 Displacement field
The Lame parameters $\lambda$ and $\mu$ depend on the density, but in this paper, we assume that the two parameters are equal constants and their values are $10^{4}\,Gpa$. The circular frequency is equal to $\omega=10$, and incident waves are made according to the formula 2.
@@ 8 Results
By dividing the dataset into training and test, each comprising 27000 and 1000 samples, respectively, the numerical and visual results of the proposed methods are given for both sets. Table 1 presents the accuracy of various methods to estimate the density function with the performance of auxiliary networks. In table 1, $\Omega$ represents the domain containing the points with the given coordinates, and $K$ denotes a set of scattered waves from each incident wave. It is important to note that since the values of $\boldsymbol{u}$ are complex numbers, their real and imaginary parts were trained separately, and the accuracy of the absolute value of them are presented in Table 1. Moreover, the parameter of $L_{2}$ regularization in the Least Squares method equals 0.02. Table 2 exhibits the values of loss functions derived from relation 21 in the third method when the parameters have been trained. Notably, the relative L2 error formulas in Table 1 and Table 2 actually represent the average of relative L2 errors calculated over the entire training and test datasets. Additionally, the value of the loss function mentioned in Algorithm 1 equals 15.
@@
Third Method (New-VAE) - - - - 0.023 0.028
@@
New-VAE 286 325 85 85 168 160 84 85 0.023 0.028
In the New-VAE row, in Tables 1 and 2, the term $\dfrac{||\rho-\overline{\rho}||_{L^{2}(\Omega)}}{||\rho||_{L^{2}(\Omega)}}$ represents the relative error between the true density $\rho$ and the reconstructed density $\overline{\rho}$. It is important to mention that this reconstructed density, as shown in position 2 of Figure 5, has been derived from the  the Linear-to-Nonlinear method. Moreover, the values of $\alpha_{1}$ and $\alpha_{2}$ used to train the third method in this paper are $0.0004$ and $0.000625$, respectively.
@@
You can visually compare the methods by looking at Figures 11 and 12, which display the results on the test dataset. In Figure 11, there is a comparison between the ground truth and deterministic techniques (first and second methods). In Figure 12, the performance of the third method is assessed against the ground truth and the second method, which provides the most accurate estimation. It is important to note that in Figure 12, the second column represents the deterministic solution of the third method (step 2 in Figure 5), equivalent to the Linear-to-Nonlinear method. The realizations of the probabilistic solution to the problem are depicted in the third, fourth, and fifth columns (step 4 in Figure 5). In particular, the third column represents the mean value of this probabilistic solution. Additionally, Figure 13 illustrates the samples produced by the generator mentioned in the third method.
Figure 11: Comparing the results of different methods with the true density
Figure 12: Some realizations from the solution of the third method compared with true density and the Linear-to-Nonlinear method
@@ 9 Discussion
In this paper, we present and compare methods based on data-driven techniques for the time-harmonic elastic FWI problem. Our study consists of several methods, including pure data-driven technique (deep learning based), the integrated application of deep learning and the physics underlying the problem, and a probabilistic deep learning technique that can also utilize the physics of the problem. Each of these methods possesses its own strengths and weaknesses. The aim of our comparative analysis is to offer guidance for researchers and experts in the field, including data scientists and geophysicists, helping them determine the most suitable method based on their specific requirements and limitations.
@@
In the following, we compare the methods presented in this paper on a case-by-case basis, according to results obtained from Table 1, Table 2, Figures  11 and 12.
@@
    In the second method, using deep learning and the physics of the problem (Lippman integral in the theory of elastic scattering), three techniques are presented to solve elastic FWI. These techniques offer advantages over PINN as they do not require auto differentiation, which itself causes errors. Additionally, since PINNs are meshless, they are incapable of finding details and have excessive smoothness. However, they are more computationally optimal than all the methods presented in this paper.
    • In the Linear-to-Nonlinear technique within this method, the simultaneous utilization of physics and machine learning demonstrates its superiority in the visual aspect. Numerically, this technique is more accurate than the first method and is resistant to noise.
@@
    In the New-VAE technique, which is a probabilistic method, we configure its deterministic component as a transcription of the Linear-to-Nonlinear method, resulting in equivalent accuracy. Moreover, it has the advantage of presenting additional solutions that might be better suited than the primary one. This capability allows experts to compare these solutions with those obtained from alternative methods and make decisions with more certainty. Additionally, this method can be used in Dataset Augmentation, as described in 5.1.
Extending the presented methods from time-harmonic to the general time-domain state can be a potential way for future exploration. Furthermore, these methods can be compared with other physics-based [22] and data-driven techniques [71], and can also be merged with them.
@@
The data generated in this paper is openly available in Mendeley Data at: Negahdari, Vahid (2024), “Elastic Full-Waveform-Inversion”, Mendeley Data, V1, doi: 10.17632/z2n2f23pxw.1.

### v2

@@ 5 Third method (New Variational Autoencoder)
We removed constants that are not relevant. Furthermore, |.||.|| . | denotes the matrix’s determinant. Given these assumptions, the generic VAE objective from 18 can be refined to:
@@ 5.1 New-VAE structue
6 Fourth method ( Physics-Based )
The physics-based method involves minimizing the distance between the data predicted by the forward model and the recorded data over the parameter space. We formulate it as the minimization problem:
$\begin{split}&\min_{\rho}\mathcal{J}\big{[}U(\rho)\big{]}=\frac{1}{2}\sum_{j=1}^{k}\left\|RU_{j}-d_{j}^{obs}\right\|^{2},\\ \text{Subject to}\hskip 14.22636pt&(U_{j}-U_{j}^{inc})_{2n^{2}\times 1}=A_{2n^{2}\times 2n^{2}}\,(\hat{\rho}U_{j})_{2n^{2}\times 1},\hskip 14.22636ptj=1,...,k\end{split}$ (23)
Where $A$, $U_{j}$, $U^{\text{inc}}_{j}$, and $\hat{\rho}$ are defined as in relation 7. The term $d_{j}^{\text{obs}}$ represents the displacement field observed on the surface. Additionally, $R$ is the mapping that projects the displacement fields $U_{j}$ in the domain $\Omega$ to their surface locations. In this paper, we address the optimization problem 23 using two approaches: the Gradient Descent and the Truncated Newton method (with Gauss Approximation). We utilize the adjoint state technique to achieve this, as described in [22].
We change the variables to make the complex-valued optimization problem 23 to a real-valued problem, which simplifies the problem.
$U_{j}=\begin{bmatrix}\Re(U_{j})\\ \Im(U_{j})\end{bmatrix}_{4n^{2}\times 1},\,\,\,U_{j}^{inc}=\begin{bmatrix}\Re(U_{j}^{inc})\\ \Im(U_{j}^{inc})\end{bmatrix}_{4n^{2}\times 1},\,\,\,A=\begin{bmatrix}\Re(A)\,\,\,\,&-\Im(A)\\ \Im(A)\,\,\,\,&\Re(A)\end{bmatrix}_{4n^{2}\times 4n^{2}},\,\,\,\rho=\begin{bmatrix}\hat{\rho},\hat{\rho}\end{bmatrix}_{1\times 4n^{2}}$ (24)
Now, with the new definitions, optimization problem 23 is defined as follows:
$\min_{\rho}\mathcal{J}\big{[}U(\rho)\big{]}=\frac{1}{2}\sum_{j=1}^{k}\left\|RU_{j}-d_{j}^{obs}\right\|^{2},\hskip 42.67912pt$ (25a)
$\text{Subject to}\hskip 14.22636pt\underbrace{(A.\rho-I)}_{A_{\rho}}U_{j}=-U_{j}^{inc}\hskip 14.22636ptj=1,...,k$ (25b)
where ”.” represents the elementwise multiplication of the vector $\rho$ in the rows of the matrix $A$.
6.1 Gradient computation
To compute the $\nabla\mathcal{J}(\rho)$ in relation 25a and 25b, it is necessary to calculate $\frac{\partial U_{j}}{\partial\rho}$, which is computationally challenging. Therefore, we use the adjoint state method to obtain the gradient. For this purpose, we define a Lagrangian function. The Lagrangian function associated with the problem 25a and 25b is given by:
$\mathcal{L}\big{[}\rho,U,\lambda\big{]}=\frac{1}{2}\sum_{j=1}^{k}\left\|RU_{j}-d_{j}^{inc}\right\|^{2}+\sum_{j=1}^{k}<A_{\rho}U_{j}+U_{j}^{inc},\lambda_{j}>$ (26)
where <.,.><.\,,.>< . , . > denotes the inner product and $\lambda_{j}$ represents the adjoint vector.
Note that $\mathcal{L}\big{[}\rho,U(\rho),\lambda\big{]}=\mathcal{J}\big{[}U(\rho)\big{]}$ independent of $\lambda$ (adjoint vector), where $U=U(\rho)$ is the forward problem solution. Differentiating this expression will yield the desired gradient.
$\nabla\mathcal{J}(\rho)=\dfrac{d}{d\rho}\mathcal{J}\big{[}U(\rho)\big{]}=\sum_{j=1}^{k}\dfrac{\ \partial\mathcal{L}}{\partial U_{j}}\dfrac{\partial U_{j}}{\partial\rho}+\frac{\partial\mathcal{L}}{\partial\rho}$ (27)
The Partial derivatives of $\mathcal{L}$ are computed as:
$\dfrac{\partial\mathcal{L}}{\partial\rho}=\sum_{j=1}^{k}(A.U_{j})^{T}\lambda_{j}$ (28a)
$\dfrac{\partial\mathcal{L}}{\partial U_{j}}=R^{T}(RU_{j}-d_{j}^{obs})+A_{\rho}^{T}\lambda_{j}\qquad j=1,...,k$ (28b)
$\dfrac{\partial\mathcal{L}}{\partial\lambda_{j}}=A_{\rho}U_{j}+U_{j}^{inc}\qquad j=1,...,k$ (28c)
To compute the $\nabla\mathcal{J}(\rho)$ as specified in relation 27, we set $\lambda_{j}$ so that $\frac{\partial\mathcal{L}}{\partial U_{j}}$ equal to zero. The gradient can then be calculated using $\frac{\partial\mathcal{L}}{\partial\rho}$. So, this process involves three steps. First, we solve the direct problem 25b to obtain $U_{j}$ at point $\rho$. In the second step, with $U_{j}$ and $\rho$ known, by setting equation 28b equals zero, we obtain the values of $\lambda_{j}$. In the third step, the $\nabla\mathcal{J}(\rho)$ is computed from equation 28a. Therefore, the Gradient Descent algorithm for problem 25a and 25b can be implemented as described in Algorithm 2.
Algorithm 2 ( Gradient Descent Algorithm )
Input : $\rho_{0},\epsilon_{1}$
      Output : $\arg\displaystyle\min_{\rho}\mathcal{J}\big{[}U(\rho)\big{]}$
1:while $\epsilon_{1}\leq\mathcal{J}\big{[}U(\rho)\big{]}$ do
2:     Solve $A_{\rho}(U_{j})=-U_{j}^{inc}\qquad j=1,...,k$
3:     Solve $A_{\rho}^{T}(\lambda_{j})=-R^{T}(RU_{j}-d_{j}^{obs})\qquad j=1,...,k$
4:     Compute $\nabla\mathcal{J}(\rho)=\sum_{j=1}^{K}(A.U_{j})^{T}\lambda_{j}$
5:end while
6:Find $\gamma$ such that it satisfies the strong Wolfe conditions
7:$\rho=\rho-\gamma\nabla\mathcal{J}(\rho)$
8: return $\rho$
Remark 6.1.
Note that the gradient vector $\nabla\mathcal{J}(\rho)$ has a length of $4n^{2}$, which results from its definition in formula 24. To ensure it is well-defined and correct, after computation of $\nabla\mathcal{J}(\rho)$ in step 4 of the Algorithm 2, it should be altered in the form of $\nabla\mathcal{J}(\rho)=\nabla\mathcal{J}(\rho)[0:n^{2}]+\nabla\mathcal{J}(\rho)[n^{2}:2n^{2}]+\nabla\mathcal{J}(\rho)[2n^{2}:3n^{2}]+\nabla\mathcal{J}(\rho)[3n^{2}:4n^{2}]$ To simplify the notation we will not mention this point in the entire of this section.
6.2 The Gauss-Newton method for approximation Hessian
Nonlinear minimization problems are typically solved using Newton-based methods. The Newton method involves building a sequence $\rho_{t}$ starting from an initial guess $\rho_{0}$, using the recurrence relation:
$\rho_{t+1}=\rho_{t}+\gamma_{t}\Delta\rho$ (29)
where $\gamma_{t}$ is a scalar parameter associated with a globalization method (linesearch) and the increment $\Delta\rho$ satisfies:
$H(\rho_{t})\Delta\rho=-\nabla\mathcal{J}(\rho_{t})$ (30)
From the definition of $\mathcal{J}$, the Hessian operator $H(\rho)$ can be written as:
$H(\rho)\quad=\quad\underbrace{\sum_{j=1}^{K}\dfrac{\partial U_{j}}{\partial\rho}^{T}R^{T}R\dfrac{\partial U_{j}}{\partial\rho}}_{H_{1}(\rho)}\quad+\quad\underbrace{\sum_{j=1}^{K}R^{T}(RU_{j}-d_{j}^{obs})\dfrac{\partial^{2}U_{j}}{\partial^{2}\rho}}_{H_{2}(\rho)}$ (31)
The matrix $H_{1}(\rho)$ is known as the Gauss-Newton approximation of the Hessian operator ,which is obtained by neglecting the $H_{2}(\rho)$ matrix.
6.3 Truncated Newton Method (with Gauss Approximation)
Unlike quasi-Newton algorithms, this method does not rely on an approximation of the inverse Hessian operator. Instead, the descent direction $\Delta\rho$ is computed as the solution to equation 30 using an iterative algorithm such as the conjugate gradient algorithm. It is important to note that in this paper, instead of considering equation 30, we use its Gauss-Newton approximation and implement the Truncated Newton method on it as follows:
$H_{1}(\rho_{t})\Delta\rho=-\nabla\mathcal{J}(\rho_{t})$ (32)
For this purpose, we consider the constrained minimization problem for an arbitrary $\{w_{j}\in\mathbb{R}^{4n^{2}}\,\,,\,\,j=1,...,k\}$:
[…]
The Lagrangian function associated with the problem 33 is:
$\mathcal{L}_{w}\big{[}\rho,U,\mu\big{]}=\sum_{j=1}^{k}<U_{j},w_{j}>+\sum_{j=1}^{k}<A_{\rho}U_{j}+U_{j}^{inc},\mu_{j}>$ (34)
Note that $\mathcal{L}_{w}\big{[}\rho,U(\rho),\mu\big{]}=\mathcal{Q}\big{[}U(\rho)\big{]}$ independent of $\mu$ (adjoint vector), where $U=U(\rho)$ is the forward problem solution. Differentiating this expression will yield the desired gradient:
$\nabla\mathcal{Q}(\rho)=\dfrac{d}{d\rho}\mathcal{Q}\big{[}U(\rho)\big{]}=\sum_{j=1}^{k}\dfrac{\ \partial\mathcal{L}_{w}}{\partial U_{j}}\dfrac{\partial U_{j}}{\partial\rho}+\frac{\partial\mathcal{L}_{w}}{\partial\rho}$ (35)
The Partial derivatives of $\mathcal{L}_{w}$ are computed as:
$\dfrac{\partial\mathcal{L}_{w}}{\partial\rho}=\sum_{j=1}^{k}(A.U_{j})^{T}\mu_{j}$ (36a)
$\dfrac{\partial\mathcal{L}_{w}}{\partial U_{j}}=A_{\rho}^{T}\mu_{j}+w_{j}\qquad j=1,...,k$ (36b)
$\dfrac{\partial\mathcal{L}_{w}}{\partial\mu_{j}}=A_{\rho}U_{j}+U_{j}^{inc}\qquad j=1,...,k$ (36c)
Note that according to 33, the relation $\nabla\mathcal{Q}(\rho)=\sum_{j=1}^{k}{\dfrac{\partial U_{j}}{\partial\rho}}^{T}w_{j}$ is resulted. By replacing $w_{j}$ with $R^{T}R\dfrac{\partial U_{j}}{\partial\rho}v$:
$\nabla\mathcal{Q}(\rho)=\sum_{j=1}^{k}{\dfrac{\partial U_{j}}{\partial\rho}}^{T}w_{j}=\sum_{j=1}^{k}\dfrac{\partial U_{j}}{\partial\rho}^{T}R^{T}R\dfrac{\partial U_{j}}{\partial\rho}v=H_{1}(\rho)v$ (37)
Therefore, it is sufficient to compute the $\dfrac{\partial U_{j}}{\partial\rho}v$ to obtain the $H_{1}(\rho)v$. To calculate $\dfrac{\partial U_{j}}{\partial\rho}v$, we take the Gateaux derivative in the direction of $v$ from equation 25b:
$A_{\rho}\Big{(}\dfrac{\partial U_{j}}{\partial\rho}v\Big{)}=-(A.U_{j})v\qquad j=1,...,k$ (38)
As a result, by solving equation 38, we obtain $w_{j}$. Then, by setting $\mu_{j}$ such that $\frac{\partial\mathcal{L}_{w}}{\partial U_{j}}=0$ , $\nabla\mathcal{Q}(\rho)$ can be calculated by $\frac{\partial\mathcal{L}_{w}}{\partial\rho}$. With $H_{1}(\rho)v$ we can now obtain the desired path $\Delta(\rho)$ by solving equation $H_{1}(\rho)v=-\nabla\mathcal{J}(\rho)$ using the conjugate gradient method. The general algorithm for this method is outlined in Algorithm 3.
Algorithm 3 ( Truncated Newton Algorithm with Gauss Approximation )
Input : $\rho_{0},\epsilon_{1},\epsilon_{2}$
      Output : $\arg\displaystyle\min_{\rho}\mathcal{J}\big{[}U(\rho)\big{]}$
1:while $\epsilon_{1}\leq\mathcal{J}\big{[}U(\rho)\big{]}$ do
2:     Solve $A_{\rho}(U_{j})=-U_{j}^{inc}\qquad j=1,...,k$
3:     Solve $A_{\rho}^{T}(\lambda_{j})=-R^{T}(RU_{j}-d_{j}^{obs})\qquad j=1,...,k$
4:     Compute $\nabla\mathcal{J}(\rho)=\sum_{j=1}^{k}(A.U_{j})^{T}\lambda_{j}$
5:     $\Delta\rho=0\,\,\,,\,\,\,r_{0}=\nabla\mathcal{J}(\rho)\,\,\,,\,\,\,v_{0}=-\nabla\mathcal{J}(\rho)\,\,\,,\,\,\,i=0$
6:     while $\epsilon_{2}\leq||r_{i}||^{2}$ do
7:         Solve $A_{\rho}\Big{(}\frac{\partial U_{j}}{\partial\rho}v_{i}\Big{)}=-(A.U_{j})v_{i}\qquad j=1,...,k$
8:         Solve $A_{\rho}^{T}(\mu_{j})=-R^{T}R\frac{\partial U_{j}}{\partial\rho}v_{i}\qquad j=1,...,k$
9:         Compute $H_{1}(\rho)v_{i}=\sum_{j=1}^{k}(A.U_{j})^{T}\mu_{j}$
10:         $\alpha_{i}=<r_{i},r_{i}>/<H_{1}(\rho)v_{i},v_{i}>$
11:         $\Delta\rho=\Delta\rho+\alpha_{i}v_{i}$
12:         $r_{i+1}=r_{i}+\alpha_{i}H_{1}(\rho)v_{i}$
13:         $v_{i+1}=-r_{i+1}+(<r_{i+1},r_{i+1}>/<r_{i},r_{i}>)v_{i}$
14:         $i=i+1$
15:     end while
16:     Find $\gamma$ such that it satisfies the strong Wolfe conditions
17:     $\rho=\rho+\gamma\Delta\rho$
18:end while
19:return $\rho$
@@ 7.1 Generating random field
To produce random fields for the dataset, we set domain $\Omega=[-1,1]\times[-1,1]$ and divide $\Omega$ to $N=n\times n$ uniform rectangular elements. Using observation and experimentation, we create 28 pairs of $(a_{1},a_{2})$ values in such a way that they cover nearly all situations of earth density. For each pair of $(a_{1},a_{2})$ we have a covariance matrix as 39, and construct 1000 realizations from it. In Figure 8, realizations obtained from specific pairs of $(a_{1},a_{2})$ are observable.
@@ 7.2 Displacement field
The Lame parameters $\lambda$ and $\mu$ depend on the density, but in this paper, we assume that the two parameters are equal constants and their values are $10^{3}\,Gpa$. The circular frequency is equal to $\omega=10$, and incident waves are made according to the formula 2.
@@ 8 Computational cost
  4. 4.
    In the fourth approach, physics-based techniques, we should solve two linear systems (the direct problem) for the main computation of gradient. This process has a computational cost of $\mathcal{O}(2kN^{3})$.
    • In the Gradient Descent technique, If we denote the number of iterations by $t_{1}$, the computational cost is $\mathcal{O}(2t_{1}kN^{3})$.
    • In the Truncated Newton (Gauss Approximation) technique, let $t_{2}$ represent the number of iterations of the inner loop in Algorithm 3. Since two linear systems are solved in each iteration of this loop, the computational cost is $\mathcal{O}(2t_{1}kN^{3}+2t_{1}t_{2}kN^{3})$.
@@ 9 Results
By dividing the dataset into training and test, each comprising 27000 and 1000 samples, respectively, the numerical and visual results of the proposed methods are given for both sets. Table 1 presents the accuracy of data-driven methods to estimate the density function with the performance of auxiliary networks. Table 2 also shows the accuracy of the physics-based method with two different initial guesses: the first guess is a constant function with a value equal to the mean of the original density, and the second initial guess is nearly close to the grand truth. In Table 1 and 2, $\Omega$ represents the domain containing the points with the given coordinates, $\partial\Omega$ refers to the surface location and $K$ denotes a set of scattered waves from each incident wave. It is important to note that since the values of $\boldsymbol{u}$ are complex numbers, their real and imaginary parts were trained separately, and the accuracy of the absolute value of them are presented in Table 1 and 2. Moreover, the parameter of $L_{2}$ regularization in the Least Squares method equals 0.02. Table 3 exhibits the values of loss functions derived from relation 22 in the third method when the parameters have been trained. Notably, the relative $L_{2}$ error formulas in Table 1, Table 2 and Table 3 actually represent the average of relative $L_{2}$ errors calculated over the entire training and test datasets. Additionally, the value of the loss function mentioned in Algorithm 1 equals 15.
@@
Third Method (New-VAE) - - - - 0.026 0.030
Table 2: Errors presented in the fourth method
$\boldsymbol{\dfrac{||u-\overline{u}||_{L^{2}(\partial\Omega\times K)}}{||u||_{L^{2}(\partial\Omega\times K)}}}$ $\boldsymbol{\dfrac{||\rho-\overline{\rho}||_{L^{2}(\Omega)}}{||\rho||_{L^{2}(\Omega)}}}$ $\boldsymbol{||\nabla\mathcal{J}(\rho)||_{L^{2}(\Omega)}}$
Initial 1 Initial 2 Initial 1 Initial 2 Initial 1 Initial 2
Gradient Descent 0.011 0.004 0.060 0.05 0.005 0.002
Physics-Based
Truncated Gauss-Newton 0.00047 0.00015 0.062 0.042 0.0002 0.00007
@@
New-VAE 55 90 81 81 50 60 83 83 0.026 0.030
In the New-VAE row, in Tables 1 and 3, the term $\dfrac{||\rho-\overline{\rho}||_{L^{2}(\Omega)}}{||\rho||_{L^{2}(\Omega)}}$ represents the relative error between the true density $\rho$ and mean of this probabilistic method which we denoted by $\overline{\rho}$. Notably, this mean is derived from samples generated from position 4 of Figure 5, using the Linear-to-Nonlinear results at position 2 of this Figure. Moreover, the values of $\alpha_{1}$ and $\alpha_{2}$ used to train the third method in this study are $1$ and $0.2$, respectively.
@@
You can visually compare the methods in Figures 11 and 12, which display the results on the test dataset. Figure 11 presents a comparison between the ground truth and data-driven methods, while Figure 12, compares the ground truth with physics-based methods, with two different initial guesses. To evaluate the third method, the probabilistic technique New-VAE, Figure 13 is provided, showing the mean and pointwise standard deviation of the VAE method alongside the ground truth and the Linear-to-Nonlinear method. The Linear-to-Nonlinear method is included because its results is used in the optimization of the New-VAE method (step 2 in Figure 5) and it achieved the best performance among the methods presented. The mean and pointwise standard deviation is calculated using the Monte Carlo method, with 5,000 samples generated in step 4 of Figure 5. Also, Figure 14 displays pointwise normalized histograms at four locations, along with the true values at those points for several density fields. Additionally, Figure 15 illustrates the samples produced by the generator mentioned in the third method.
Figure 11: Comparison of data-driven methods’ results with the true density
Figure 12: Comparison of physics-based methods’ results, using two different initial guesses, with the true density
Figure 13: Mean and pointwise standard deviation of the third method, compared with true density and the results from the Linear-to-Nonlinear method
Figure 14: Pointwise normalized histogram at four locations, shown alongside the true values at these points for some density fields
@@ 10 Discussion
In this paper, we present and compare methods based on physics-based and data-driven techniques for the time-harmonic elastic FWI problem. Our study consists of several methods, including physics-based, pure data-driven technique (deep learning based), the integrated application of machine learning and the physics underlying the problem, and a probabilistic deep learning technique that can also utilize the physics of the problem. Each of these methods possesses its own strengths and weaknesses. The aim of our comparative analysis is to offer guidance for researchers and experts in the field, including data scientists and geophysicists, helping them determine the most suitable method based on their specific requirements and limitations.
@@
In the following, we compare the methods presented in this paper on a case-by-case basis, according to results obtained from Tables 1, 2, 3, and Figures  11, 12, 13 and 14.
@@
    In the second method, using machine learning and the physics of the problem (Lippman integral in the theory of elastic scattering), three techniques are presented to solve elastic FWI. These techniques offer advantages over PINN as they do not require auto differentiation, which itself causes errors. Additionally, since PINNs are meshless, they are incapable of finding details and have excessive smoothness. However, they are more computationally optimal than all the methods presented in this paper and doesnt need to dataset.
    • In the Linear-to-Nonlinear technique within this method, the simultaneous utilization of physics and machine learning demonstrates its superiority in the visual aspect. Numerically, this technique is more accurate than other deterministic methods presented in this paper.
@@
    In the New-VAE technique, which is a probabilistic method, we configure its deterministic component as a transcription of the Linear-to-Nonlinear method. By visually comparing the mean of this method with that of other approaches, it reveals that this method demonstrates resistance to noise. Moreover, it offers the advantage of performing uncertainty quantification analysis and proposing alternative solutions that may be more suitable than the primary one. This capability allows experts to compare these solutions with those obtained from alternative methods and make decisions with more certainty. Additionally, this method can be used in Dataset Augmentation, as described in 5.1.
  4. 4.
    In the fourth method, Physics-Based approach, which is included for comparison with data-driven methods, we employ two techniques: Gradient Descent and Truncated Gauss Newton method. For each technique, we use two different initial guesses, as shown in Table 2 and Figure 12. These results indicate that when the initial guess is close to the ground truth, the accuracy of these methods, particularly the Gauss method, improves significantly. This indicates that a suitable initial guess is essential for achieving high accuracy, which represents a limitation of this approach since a suitable initial guess may not always be available. Although the accuracy of this physics-based method is lower than that of methods like Linear-to-Nonlinear, it is expected that with better initial guesses, its accuracy can improve.
Extending the presented methods from time-harmonic to the general time-domain state can be a potential way for future exploration. Moreover, obtaining additional parameters in elastic FWI, such as P-wave and S-wave velocities, requires the use of a Green’s function for scenarios involving non-constant Lame parameters. This Green’s function may be derived analytically or approximated numerically, offering a promising direction for future research. Furthermore, the presented methods in this paper can be compared with other physics-based [22] and data-driven techniques [77], and can also be merged with them.
@@
The dataset generated in this paper is openly available in Mendeley Data at [101].